# Leveraging Edge Intelligence and LLMs to Advance 6G-Enabled Internet of Automated Defense Vehicles

### v1

@@ I Introduction
Automated defense Vehicles (ADVs) in military settings offer strategic benefits, including reducing risk to soldiers, enhancing reconnaissance, and improving logistics and supply chain efficiency in combat zones. These defense vehicles must navigate and operate in hazardous, complex environments with minimal human intervention. However, challenges arise in military settings, such as terrain awareness, off-road navigation, unfamiliar environments, the need for complete re-routing in open areas, identifying alternate routes, and ensuring optimal vehicle control [2]. To address these challenges, integrating multiple connected vehicles into the 6G-enabled Internet of Automated Defense Vehicles (IoADV) framework can significantly improve situational awareness, coordination, and operational efficiency [3]. The 6G-enabled IoADV allows real-time data sharing, autonomous convoy operations, and synchronized maneuvers, ensuring safe and efficient missions in hostile environments. Additionally, military-grade AI faces ethical and legal challenges, including transparency in decision-making and liability for errors or collateral damage. Thus managing large volume data, ensuring real-time processing, maintaining reliable communication, efficient decision-making, and addressing cybersecurity threats are critical issues [1].
Specifically, there are two main challenges: reliable communication and decision-making for autonomous driving of ADVs. To enhance communication, the use of a large generative AI model is proposed for 6G-enabled IoADV. This model leverages advanced data processing, multi-modal integration, and predictive analytics to improve communication and coordination among vehicles [4]. This approach aims to enable more reliable, efficient IoADVs by utilizing generative and predictive capabilities of AI models to optimize various network functions and support real-time decision-making. For decision-making in autonomous driving, a framework that integrates pre-trained generative Multimodal Large Language Models (M-LLMs) [5, 6] is proposed due to their advanced features to assist ADVs by integrating and processing diverse data types such as visual inputs, sensor readings, and textual information. This integration enhances vehicle’s ability to understand and respond to complex driving environments, improving perception, decision-making, and real-time navigation. Moreover, LLMs provide transparency and explainability in decision-making, addressing the “black box” issue associated with deep neural networks and other advanced AI models. To the best of our knowledge, this is the first time in the literature that LLMs are being explored within the context of 6G-enabled IoADVs for military applications.
@@ II-B Requirement of Military Defense Services
Military defense requirements encompass a wide range of technological and ethical considerations to ensure effective and secure operations. Though autonomous vehicle research is common in the civilian domain, it can not be directly integrated into the military domain. One of the reasons is civilian autonomous vehicles often lack the robust security measures required for military operations, making them susceptible to cyberattacks and electronic warfare. Adversaries can exploit unencrypted communication links to intercept data or disrupt operations through jamming and spoofing attacks. The reliance on commercial off-the-shelf technologies without adequate hardening can compromise mission integrity.
Moreover, civilian technologies are typically designed for controlled environments and may not withstand the harsh conditions of military operations, such as extreme temperatures, dust, and electromagnetic interference. Also, integrating civilian systems with existing military infrastructure poses significant interoperability issues. Differences in communication protocols, data formats, and operational standards can hinder seamless integration, leading to inefficiencies and potential mission failures. Ensuring civilian technologies can operate cohesively with military systems requires extensive modifications and rigorous testing [7].
@@ III-A Enhanced Communication in IoADV via 6G and Edge Computing
Moreover, the integration of advanced AI and edge computing enables sophisticated data processing and decision-making at the network’s edge, significantly reducing latency and response times. Edge computing improves AI functionality in ADV vehicles by positioning computational resources closer to the data source, ensuring real-time processing of sensor data critical for immediate decision-making in dynamic environments. Our proposed solution utilizes edge intelligence [9] for connected vehicles, offering scalability while lowering costs and supporting URLLC and real-time data exchange to satisfy strict, seamless communication requirements of AI models on ADV in military domains. In military applications, ADVs gather multimodal sensor data, which is transmitted to edge devices via 6G, enabling the IoADV. The edge unit processes the data and returns outputs to the vehicles for immediate action. The workflow of the proposed solution is illustrated in Figure3.
@@ III-B Security Concern of Communication
The increased connectivity and multi-layered architecture of IoT expose systems to cybersecurity risks, where a breach in one device can compromise the entire network. In military contexts, compromised IoADV devices risk unauthorized access to sensitive data, manipulation, and communication disruption. These devices face unique challenges due to their critical nature and hostile environments [10]. Strong security measures are essential to protect sensitive information and maintain operational integrity. There are several suggestions to overcome these security challenges in IoT communication throughout the literature. To enhance trust and resilience, military communication systems can adopt distributed learning models and cryptographic techniques, ensuring adaptability and reliability. Advanced network protocols like semantic edge-based systems improve scalability and interoperability, while lightweight access control protocols secure communication through multi-factor authentication and session key establishment. Threat modeling techniques, including STRIDE and DREAD, enable systematic identification and mitigation of vulnerabilities. Finally, decision control frameworks, such as emotion-aware systems and multi-level data fusion, enhance battlefield situational awareness and support real-time, precise operational decisions [10].
@@ IV-A Network Optimization Using LLM and Generative AI
LLMs and Generative AI (GenAI) significantly impact the field of telecommunications by enhancing wireless communication and enabling more intelligent, autonomous networks. These models improve situational and contextual awareness by leveraging multimodal data, such as radio frequency (RF) signals and 2D/3D representations of wireless environments. Their predictive capabilities allow networks to proactively manage tasks like localization, beamforming, power allocation, and spectrum management. Another benefit of GenAI is to excel in cyber threats such as prompt injection, insecure output handling, data poisoning, DDoS attacks, jamming, and so on.
@@
Large-GenAI models rely on massive raw data, leading to significant computational overhead and redundancy. Semantic communication, which focuses on transmitting the meaning and intent behind data rather than raw data itself, can help reduce transmission size and improve network throughput. This approach requires redesigning the wireless physical layer to support efficient, semantic-based communication, allowing large GenAI models to learn from compressed and abstracted data represented as knowledge base. By understanding the intent and meaning behind data, semantic communication can optimize network resources by prioritizing critical and contextually important information, thereby improving communication quality and resource utilization in 6G environment. For example, during a joint military operation, semantic communication can prioritize transmitting mission-critical information, such as enemy positions or troops movements, while abstracting less critical data. This not only enhances resource utilization but also secures communication channels by focusing on transmitting knowledge rather than raw data.
Emergent protocol learning in generative AI enables AI agents to autonomously develop and adapt communication protocols via interactions, often using Multi-Agent Reinforcement Learning (MARL). This fosters goal-driven, compositional, and context-aware wireless MAC and network protocols, optimizing resource allocation and improving ADV control. Large GenAI models enhance MARL by guiding collaborative actions with expert planning, reducing data requirements and improving convergence. Distributed GenAI models further enable real-world task planning, creating coordinated action sequences and delegating sub-tasks to devices, enhancing efficiency through cooperation.
@@
For instance, during coordinated military tank strike, Multi-agent reinforcement learning (MARL)-trained agents can dynamically adjust protocols to optimize resource allocation for navigation, target acquisition, and secure data exchange. Large GenAI models act as guides for MARL, offering expert planning capabilities that improve convergence and reduce the need for extensive data collection. Another example of these capabilities in action is distributed military reconnaissance mission involving drones and ground units. In such scenario, distributed large-GenAI models can create action sequences for the collaborative vehicles, assigning tasks like mapping enemy territories, identifying threats, and transmitting findings to command units. By leveraging real-world grounding and cooperative planning, they can operate autonomously, sharing critical insights while minimizing redundancy. This approach enhances mission execution efficiency, improves resource utilization, and ensures seamless coordination between all units involved in the operation.
On the other hand, LLMs play a pivotal role in handling cybersecurity threats by leveraging their advanced natural language understanding, pattern recognition, and decision-making capabilities. One of their primary strengths lies in threat detection and analysis. LLMs analyze vast amounts of network data collected in IoADV sensors in real-time, identifying anomalies and recognizing patterns indicative of cybersecurity threats, such as malware, jamming signals, and unusual traffic behaviors. Their ability to process and contextualize large datasets allows them to flag potential issues before they escalate, outperforming traditional models in areas like malware detection and intrusion detection [12].
In a military operation, command and control systems are the backbone of effective decision-making and coordination, often functioning in dynamic and high-risk environments. These systems are vulnerable to various cyber threats, including Advanced Persistent Threats (APTs), phishing campaigns, jamming, and DDoS attacks. LLMs can play a crucial role in securing these systems by providing real-time monitoring, detecting sophisticated cyber threats, and automating responses to maintain the integrity and availability of critical resources.
@@ IV-B Decision-Making in ADVs by Multimodal-LLM
Conventional autonomous driving systems rely on rule-based algorithms for route planning but struggle in unforeseen scenarios like rare accidents, surprise attacks, or unknown territory. Advanced DL models outperform these methods in planning and decision-making but face challenges due to their black-box nature, raising ethical and legal concerns [13, 6]. Issues such as model illusions can cause misinterpretations of the environment, leading to safety risks, while biases may result in unfair decisions across diverse environments. Additionally, errors in reasoning and false information can lead to inappropriate or dangerous driving behaviors. Moreover, inductive advice might expose vehicles to external interference or malicious actions [5].
Therefore, explainable autonomous driving has gained attention for its ability to clarify the “black box” nature of decision-making. Multimodal LLMs offer advanced capabilities in understanding and generating human-like language, which can enhance explainability and decision-making in autonomous driving. First, LLMs improve vehicle-person interaction by understanding and responding to natural language commands, enabling more intuitive and seamless remote communication. Additionally, LLMs enhance decision-making by integrating contextual information from sources like roads, civilians, targets, weather, unknown territory, and obstacles. Moreover, using specialized fine-tuning and adaptation methods, multimodal-LLMs can easily cover army-specific knowledge, doctrines and terminologies, mission purpose, and foreign environment during the mission to improve the model’s ability to generate contextually appropriate and accurate responses for military applications, which is also essential since LLMs are trained with mainly daily scenario, but military domain requires specific knowledge. They also assist in the continuous learning and improvement of autonomous systems by analyzing large datasets of driving scenarios, accidents, enemy territory, and feedback, providing insights for system improvement as well as ethical considerations. For instance, in situations where an autonomous vehicle must decide whether to engage a target, an LLM can evaluate the scenario against established criteria, such as distinguishing between combatants and non-combatants, assessing proportionality, and ensuring necessity, thereby supporting compliance with ethical standards. LLM-assisted ADVs can also provide proper feedback by giving prompts and enhancing the military operation’s direction. The workflow of the proposed framework is shown in Figure5.
@@
Several studies have been conducted to improve the LLM’s performance in autonomous driving, as shown in Table I. Most studies utilize images and user prompts as inputs, while GPS and radar signal data were collected in [5] and [6]. In [14], to provide continuous learning from experiences and improve further action, lifelong learning is applied. Through this learning, the autonomous driving system continuously acquires, adapts, and fine-tunes its knowledge and skills over time. The reinforcement Learning (RL) model and DriveLLM framework are adopted in [6] for Self-Reflection, where it evaluates decisions and summarizes mistakes. For training, there are two-tier RL processes: cyber and physical environmental process, where the model learns from both virtual and real-world environments and receives feedback to ensure continuous and adaptable learning, and Iterative Online Learning, where the system employs iterative online learning by leveraging the feedback from both simulations and real-world executions. It is tested on a real-world WATonoBus platform, and the DriveLLM framework achieved a median decision-making time of 0.09 seconds and operated at 5–13 decisions per second (DPS) using the GPT-3.5 Turbo backend. RL is also leveraged in [14] to enable continuous learning and improvement in ADVs by evaluating and reflecting on the agent’s decisions and providing detailed feedback to enhance future behavior. Performance metrics of this work are the effectiveness in lifelong learning and collaboration success rate. This method significantly improves success steps (SSmean), such as increasing from 8.2 to 17.6 in the Intersection scenario with the use of memory shots, compared to traditional methods. Additionally, its communication and reflection modules boosted performance, improving the success rate from 50% to 80% in collaborative scenarios. Furthermore, research [5] is performed in various datasets and simulation environments, such as CARLA, and Combines LLMs with vision-based models. Tasks include planning, perception, and reasoning with multimodal inputs such as text, images, and LiDAR. One of the datasets, such as DRAMA, stands out as a comprehensive resource with 91 hours of data, 17,785 videos, and 77,639 QA pairs, emphasizing risk localization and visual reasoning for autonomous driving. It sets a benchmark for improving decision-making and reasoning capabilities.
@@ V Opportunities and Challenges
Applying ADVs in military domains provides numerous opportunities and streamlines operations such as terrain analysis, route optimization, medical evacuation, and border security without requiring human intervention. Integrating multimodal LLMs into the decision-making processes of ADVs helps meet military demands in areas such as transparency, adaptability, confidentiality, robustness, and the ability to handle vast amounts of data.Moreover, using multiple ADVs in military operations, communicating with each other and their surroundings via 6G communication, enhances perception by gathering data from various locations and enabling cooperative work. This reduces the workload on individual AV. The integration of 6G and edge computing further improves the robustness and reliability of IoADV while reducing inference times, making it well-suited for real-time decision-making. Additionally, incorporating GenAI models optimizes the network, ensuring seamless communication.However, applying these technologies to the military domain faces some key challenges as listed below:

### v2

@@ I Introduction
Automated defense Vehicles (ADVs) in military settings offer strategic benefits, including reducing risk to soldiers, enhancing reconnaissance, and improving logistics and supply chain efficiency in combat zones. These defense vehicles must navigate and operate in hazardous, complex environments with minimal human intervention. However, challenges arise in military settings, such as terrain awareness, off-road navigation, unfamiliar environments, the need for complete re-routing in open areas, identifying alternate routes, and ensuring optimal vehicle control [2]. To address these challenges, integrating multiple connected vehicles into the 6G-enabled Internet of Automated Defense Vehicles (IoADV) framework can significantly improve situational awareness, coordination, and operational efficiency. The 6G-enabled IoADV allows real-time data sharing, autonomous convoy operations, and synchronized maneuvers, ensuring safe and efficient missions in hostile environments. Additionally, military-grade AI faces ethical and legal challenges, including transparency in decision-making and liability for errors or collateral damage. Thus managing large volume data, ensuring real-time processing, maintaining reliable communication, efficient decision-making, and addressing cybersecurity threats are critical issues [1].
Specifically, there are two main challenges: reliable communication and decision-making for autonomous driving of ADVs. To enhance communication, the use of a large generative AI (GenAI) model is proposed for 6G-enabled IoADV. This model leverages advanced data processing, multi-modal integration, and predictive analytics to improve communication and coordination among vehicles [3]. This approach aims to enable more reliable, efficient IoADVs by utilizing generative and predictive capabilities of AI models to optimize various network functions and support real-time decision-making. For decision-making in autonomous driving, a framework that integrates pre-trained generative Multimodal Large Language Models (M-LLMs) [4] is proposed due to their advanced features to assist ADVs by integrating and processing diverse data types such as visual inputs, sensor readings, and textual information. This integration enhances vehicle’s ability to understand and respond to complex driving environments, improving perception, decision-making, and real-time navigation. Moreover, LLMs provide transparency and explainability in decision-making, addressing the “black box” issue associated with deep neural networks and other advanced AI models. To the best of our knowledge, this is the first time in the literature that LLMs are being explored within the context of 6G-enabled IoADVs for military applications.
@@ II-B Requirement of Military Defense Services
Military defense operations demand robust technological and ethical considerations, making direct integration of civilian autonomous vehicles unfeasible. These vehicles often lack the necessary security measures, leaving them vulnerable to cyberattacks, jamming, and spoofing. Additionally, their reliance on commercial off-the-shelf components compromises mission integrity due to inadequate hardening. Beyond security risks, civilian technologies are designed for controlled environments and may fail under extreme military conditions, such as harsh weather and electromagnetic interference. Interoperability challenges further complicate integration with military infrastructure, as differences in communication protocols and operational standards require extensive modifications and rigorous testing [5].
@@ III-B Security Concern of Communication
The increased connectivity and multi-layered architecture of IoT expose systems to cybersecurity risks, where a breach in one device can compromise the entire network. In military contexts, compromised IoADV devices risk unauthorized access to sensitive data, manipulation, and communication disruption. These devices face unique challenges due to their critical nature and hostile environments [8]. Robust security measures are crucial for protecting sensitive data and maintaining operational integrity. Military communication systems can enhance resilience using distributed learning models and cryptographic techniques. Semantic edge-based systems improve scalability, while lightweight access control protocols ensure secure communication via multi-factor authentication. Threat modeling (e.g., STRIDE, DREAD) helps identify vulnerabilities, and decision control frameworks, like emotion-aware systems and multi-level data fusion, enhance battlefield awareness and real-time decision-making [8].
@@ IV-A Network Optimization Using LLM and Generative AI
LLMs and GenAI significantly impact the field of telecommunications by enhancing wireless communication and enabling more intelligent, autonomous networks. These models improve situational and contextual awareness by leveraging multimodal data, such as radio frequency (RF) signals and 2D/3D representations of wireless environments. Their predictive capabilities allow networks to proactively manage tasks like localization, beamforming, power allocation, and spectrum management. Another benefit of GenAI is to excel in cyber threats such as prompt injection, insecure output handling, data poisoning, DDoS attacks, jamming, and so on.
@@
Large GenAI models generate massive raw data, leading to high computational overhead and redundancy. Semantic communication mitigates this by transmitting only the intent and meaning behind data, reducing transmission size and optimizing network throughput. This requires redesigning the wireless physical layer to enable knowledge-based learning from compressed, abstracted data. By prioritizing mission-critical information, such as enemy positions in military operations, semantic communication enhances resource utilization and secures transmissions. Moreover, Emergent protocol learning allows AI agents to autonomously develop adaptive communication protocols using Multi-Agent Reinforcement Learning (MARL), enabling context-aware MAC and network optimization. Large GenAI models enhance MARL by guiding collaborative actions, reducing data needs, and improving convergence. Distributed GenAI models further streamline real-world task planning by coordinating actions and delegating sub-tasks, improving efficiency through cooperative intelligence.
@@
During a coordinated military tank strike, MARL-trained agents dynamically adjust protocols for navigation, target acquisition, and secure data exchange, while large GenAI models enhance planning, improving convergence and reducing data requirements. In distributed reconnaissance missions, GenAI models coordinate drones and ground units, assigning tasks like mapping enemy territories, identifying threats, and transmitting intelligence. By leveraging cooperative planning and real-world grounding, they enable autonomous operations, optimizing resource use and ensuring seamless coordination [10].
On the other hand, LLMs are crucial in countering cybersecurity threats by leveraging natural language understanding, pattern recognition, and decision-making abilities. They excel at real-time threat detection and analysis, particularly against jamming signals that pose significant risks to ADVs, which depend on wireless communication for receiving commands and applying decision-making. While Frequency Hopping Spread Spectrum (FHSS) is a common anti-jamming technique, military environments require more adaptive solutions due to the higher stakes of jamming attacks. LLMs can enhance the security of command-and-control systems by monitoring and detecting advanced threats like Advanced Persistent Threats (APTs), phishing campaigns, jamming, and DDoS attacks, and automating responses to protect critical resources. In anti-jamming protocols, LLMs dynamically adapt by analyzing attack types and severity, optimizing frequency hopping, adjusting beamforming, and fine-tuning power levels to ensure robust and resilient communication.
@@ IV-B Decision-Making in ADVs by Multimodal-LLM
IV-B1 Advancement Over Traditional Methods
Conventional autonomous driving systems rely on rule-based algorithms for route planning but struggle in unforeseen scenarios like rare accidents, surprise attacks, or unknown territory. Advanced DL models outperform these methods in planning and decision-making but face challenges due to their black-box nature, raising ethical and legal concerns [4]. Issues such as model illusions can cause misinterpretations of the environment, leading to safety risks, while biases may result in unfair decisions across diverse environments.
Explainable autonomous driving is crucial for clarifying the “black box” nature of decision-making. Multimodal LLMs enhance explainability by processing human-like language, improving vehicle-person interaction, and integrating contextual data from roads, civilians, weather, and obstacles. For military applications, fine-tuned LLMs adapt to domain-specific knowledge, including doctrines, mission objectives, and foreign environments, addressing gaps in traditional training datasets. They also facilitate continuous learning by analyzing driving scenarios, enemy territories, and operational feedback. In critical decisions, such as target engagement, LLMs assess compliance with ethical standards, distinguishing combatants from non-combatants while ensuring proportionality and necessity. Additionally, LLM-assisted ADVs enhance military operations by providing real-time feedback and decision support. The proposed framework’s workflow is shown in Figure5.
IV-B2 Ethical Consideration of LLM
Addressing ethical concerns in military operations necessitates adherence to established frameworks and protocols that ensure compliance with international standards. A structured framework is proposed to evaluate the ethical implications of emerging technologies in military contexts in research [11]. This framework emphasizes the importance of aligning technological developments with established ethical principles and international standards, such as the Law of Armed Conflict (LOAC). By systematically assessing factors such as proportionality, distinction, and necessity, the framework aims to ensure that new technologies are integrated into military operations responsibly and ethically. The authors advocate for proactive ethical evaluations to anticipate potential dilemmas and guide the development and deployment of technologies in a manner consistent with both national and international ethical obligations.
@@ IV-B2 Ethical Consideration of LLM
IV-B3 Literature Review of Multimodal LLM and Autonomous Vehicles
Several studies have enhanced LLM performance in autonomous driving (Table I). Most use images and user prompts, while GPS and radar data were incorporated in [4]. Lifelong learning in [12] enables continuous adaptation, allowing the system to refine its knowledge over time. The DriveLLM framework in [4] employs Reinforcement Learning (RL) for self-reflection, assessing decisions, and summarizing mistakes. Its training follows a two-tier RL process, learning from both cyber and physical environments with iterative feedback. It achieved a 0.09s median decision-making time and 5–13 decisions per second (DPS) using GPT-3.5 Turbo. Similarly, RL in [12] improves continuous learning by evaluating past decisions and refining future behavior. Performance metrics include lifelong learning effectiveness and collaboration success rate, with success steps (SSmean) increasing from 8.2 to 17.6 in intersections using memory shots. Communication and reflection modules further boosted collaboration success from 50% to 80%, demonstrating adaptive learning benefits.
LLMs enhance resilience in adversarial military scenarios by mitigating risks and reducing latency. HUDSON [13] strengthens autonomous driving against perception attacks using multimodal inputs (images, RADAR/LIDAR) and LLM-based reasoning. It improves object detection, motion planning, and decision-making, ensuring robustness against adversarial manipulations. Evaluations under simulated attacks assess attack success rate, detection accuracy, decision correctness, and adversarial robustness. The performance metrics include attack success rate, object detection accuracy, decision correctness in motion planning, adversarial robustness, and the impact of perception errors on driving outcomes. Work [14] addresses long-tail driving scenarios by leveraging LLMs for behavior planning, trajectory optimization, and safety verification. Experiments show LLM-driven decision-making enhances safety, computational efficiency, and driving performance. Compared to the DriveLikeAHuman baseline, it ensures non-collision behavior in 1500 steps, improves average speed (34.3±7.7 m/s vs. 25.0±5.1 m/s), and significantly reduces latency per decision (1.7±2.7s vs. 31.1±12.2s). performance metrics include safety assurance, time efficiency in terms of computational latency measured by LLM’s response time, and driving performance metrics, including average speed. Long-tail scenarios, often rare and complex, challenge traditional models due to limited training data, but LLMs offer adaptive solutions. In military operations, vehicles mostly encounter this kind of situation where roads and buildings are deconstructed or terrible weather conditions, and adversarial attacks mostly happen during the decision-making in the operations. These works provide better decision-making under uncertainty, robustness for adaptation, and faster actions for ADV.
@@ V Opportunities and Challenges
Applying ADVs in military domains provides numerous opportunities and streamlines operations such as terrain analysis, route optimization, medical evacuation, and border security without requiring human intervention. Integrating multimodal LLMs into the decision-making processes of ADVs helps meet military demands in areas such as transparency, adaptability, confidentiality, robustness, and the ability to handle vast amounts of data. Moreover, using multiple ADVs in military operations, communicating with each other and their surroundings via 6G communication, enhances perception by gathering data from various locations and enabling cooperative work. This reduces the workload on individual AV. The integration of 6G and edge computing, where edge devices, close to the data source, leverage collaborative approaches to share their resource and distribute the computation load of given tasks, further improves the robustness, scalability, and reliability due to keeping data on edge devices, of IoADV while reducing inference times in low resource environments, making it well-suited for real-time decision-making in the military domain. [6]. Additionally, incorporating GenAI models optimizes the network, ensuring seamless communication. However, applying these technologies to the military domain faces some key challenges as listed below:
@@ V-B Training LLMs for Military Applications
However, military domains require complete verification for synthetic data to ensure that artificial datasets accurately represent real-world scenarios due to the reliability of the datasets for model training [1]. One of the well-known verification methods is called statistical analysis, which ensures alignment in distributions, correlations, and variances. This comparison confirms that the synthetic data accurately reflects the characteristics of actual operational environments. If synthetic data is collected through the simulation, it is tested in simulated military scenarios using relevant operational parameters, such as battlefield topography, weather conditions, enemy tactics, and so on, to assess its realism. Moreover, incorporating military domain expertise is crucial for the effective validation of synthetic data and identifying potential inconsistencies in defense applications.